# Diffusion on language model encodings for protein sequence generation

### v2

@@ E.2 Comparison Across Generative Paradigms on SwissProt Dataset
Distributional Similarity Dataset 0.13 0.00 1.08 0.00 0.00 0.05
@@
Diversity and Novelty Dataset 0.045 1.000 0.943 0.990 0.304 25.3
@@ E.5 Protein Family Controllable Generation
Classifier Guidance. In the classifier guidance approach, we train a classifier that takes as input a noisy, normalized encoding of ESM2-650M along with a time label and predicts logits for $9$ classes. Label 00 represents the absence of a class, where a protein is randomly sampled from the AFDB in $50\%$ of cases, ensuring it does not belong to any of the specified protein families. The remaining labels correspond to specific protein families. The classifier architecture consists of three transformer blocks with latent attention (Lee et al., 2024), followed by a linear layer. Training is performed with a batch size of $1024$ and a learning rate of $42e-$ 4 over $300$ iterations.
Conditional Fine-tuning. In the conditional fine-tuning approach, we also utilize $9$ labels and randomly sample a protein from the AFDB in $50\%$ of cases for the absence of a specific family. To condition the diffusion model on family labels, we introduce an embedding layer. The label embeddings, after passing through a linear layer, are incorporated into the input of each transformer block within the diffusion model. The fine-tuning process is conducted with a batch size of 1024 over 2500 iterations, using a learning rate of $55e-$ 5.
@@
For training ProLLAMA, we initialize the model with pre-trained weights available at https://huggingface.co/GreatCaptainNemo/ProLLaMA. Fine-tuning is performed using LoRA (Hu et al., 2021) on an annotated dataset comprising eight protein families, following the format recommended by the original authors. During our experiments, we observed that the generation parameters suggested by the authors were suboptimal for producing family-specific protein data. Most generated proteins did not correspond to the specified family. To address this issue, we adjusted the parameters and identified a set of values that yielded the highest generation fidelity. The optimized parameters are: $top_{k}=40$, $top_{p}=0.9$, repetition penalty$=1.2$, temperature$=0.5$. The most significant modification was in the temperature parameter, where we replaced the originally suggested value of $0.2$ with $0.5$, resulting in improved family-specific protein generation.
For fine-tuning ProGen2, we use the code provided in the repository https://github.com/hugohrban/ProGen2-finetuning and initialize the model with pre-trained weights from https://github.com/salesforce/progen/tree/main/progen2. To condition the model on the family label, we prepend the family tag at the beginning of each sequence. For sampling, we use the following parameters: temperature $=0.5$ and $top_{k}=40$.
@@ E.6 Joint Sequence-Structure Generation with DiMA
[…]
@@
[…]
@@ Appendix F Biological Relevance Analysis
Figure 18: Analysis of protein domain characteristics in generated sequences. (A) Fraction of sequences containing InterPro SUPERFAMILY domains shows comparable annotation rates between models and natural proteins. (B) Venn diagram revealing DiMA’s superior domain coverage, generating both more unique domains (67) and higher total domain count compared to DPLM and nanoGPT. (C) Domain length distributions demonstrating that DiMA accurately preserves natural domain length patterns, while DPLM shows bias toward full-sequence domains (240–250 aa), and nanoGPT exhibits non-native uniform distribution. These results indicate DiMA’s capacity to capture diverse and biologically relevant domain architectures.

### v4

@@ E.2 Comparison Across Generative Paradigms on SwissProt Dataset
Distributional Similarity
Dataset 0.13 0.00 1.08 0.00 0.00 0.05
@@
Diversity and Novelty
Dataset 0.045 1.000 0.943 0.990 0.304 25.3
@@ E.5 Protein Family Controllable Generation
Classifier Guidance. In the classifier guidance approach, we train a classifier that takes as input a noisy, normalized encoding of ESM2-650M along with a time label and predicts logits for $9$ classes. Label 0 represents the absence of a class, where a protein is randomly sampled from the AFDB in $50\%$ of cases, ensuring it does not belong to any of the specified protein families. The remaining labels correspond to specific protein families. The classifier architecture consists of three transformer blocks with latent attention (Lee et al., 2024), followed by a linear layer. Training is performed with a batch size of $1024$ and a learning rate of 2e42e-4 over $300$ iterations.
Conditional Fine-tuning. In the conditional fine-tuning approach, we also utilize $9$ labels and randomly sample a protein from the AFDB in $50\%$ of cases for the absence of a specific family. To condition the diffusion model on family labels, we introduce an embedding layer. The label embeddings, after passing through a linear layer, are incorporated into the input of each transformer block within the diffusion model. The fine-tuning process is conducted with a batch size of 1024 over 2500 iterations, using a learning rate of 5e55e-5.
@@
For training ProLLAMA, we initialize the model with pre-trained weights available at https://huggingface.co/GreatCaptainNemo/ProLLaMA. Fine-tuning is performed using LoRA (Hu et al., 2021) on an annotated dataset comprising eight protein families, following the format recommended by the original authors. During our experiments, we observed that the generation parameters suggested by the authors were suboptimal for producing family-specific protein data. Most generated proteins did not correspond to the specified family. To address this issue, we adjusted the parameters and identified a set of values that yielded the highest generation fidelity. The optimized parameters are: $top_{k}=40$, $top_{p}=0.9$, repetition penalty=1.2=1.2, temperature=0.5=0.5. The most significant modification was in the temperature parameter, where we replaced the originally suggested value of $0.2$ with $0.5$, resulting in improved family-specific protein generation.
For fine-tuning ProGen2, we use the code provided in the repository https://github.com/hugohrban/ProGen2-finetuning and initialize the model with pre-trained weights from https://github.com/salesforce/progen/tree/main/progen2. To condition the model on the family label, we prepend the family tag at the beginning of each sequence. For sampling, we use the following parameters: temperature =0.5=0.5 and $top_{k}=40$.
@@ E.6 Joint Sequence-Structure Generation with DiMA
A generated protein is considered designable if it satisfies the criterion scRMSD $2\mathring${A} and scTM >0.5>0.5. Following established literature (Yim et al., 2023b; Lin et al., 2024; Campbell et al., 2024; Bose et al., 2023), we also report the proportion of designed proteins that meet this self-consistency threshold.
@@
We present the distributions of self-consistency metrics in Fig. 18, demonstrating the high designability of the generated sequences and structures. The results indicate that the predicted amino acid sequences and backbone conformations are highly consistent, with $89\%$ of proteins achieving scRMSD << $2\mathring${A} and $99\%$ exhibiting scTM >> 0.5. To further assess the designability of DiMA, we compare its performance against structure-based generative models (Watson et al., 2023; Yim et al., 2023b; Lin et al., 2024; Bose et al., 2023; Ingraham et al., 2023) in Table 11. DiMA’s performance is evaluated using two complementary approaches:
@@ Appendix F Biological Relevance Analysis
Figure 23: Analysis of protein domain characteristics in generated sequences. (A) Fraction of sequences containing InterPro SUPERFAMILY domains shows comparable annotation rates between models and natural proteins. (B) Venn diagram revealing DiMA’s superior domain coverage, generating both more unique domains (67) and higher total domain count compared to DPLM and nanoGPT. (C) Domain length distributions demonstrating that DiMA accurately preserves natural domain length patterns, while DPLM shows bias toward full-sequence domains (240–250 $a$), and nanoGPT exhibits non-native uniform distribution. These results indicate DiMA’s capacity to capture diverse and biologically relevant domain architectures.